The equation for a distance, s(m), travelled in time t(s) by an object starting with an initial velocity u(ms-1) and uniform acceleration a(ms-2) is: 𝑠=𝑒𝑑+12π‘Žπ‘‘2 The tasks are to: a) Plot a graph of distance (s) vs time (t) for the first 10s of motion if 𝑒=10π‘šπ‘ βˆ’1 and π‘Ž=4π‘šπ‘ βˆ’2. b) Determine the gradient of the graph at 𝑑=2𝑠 and 𝑑=6𝑠. c) Differentiate the equation to find the functions for i) Velocity (𝑣=𝑑𝑠𝑑𝑑) ii) Acceleration (π‘Ž=𝑑𝑣𝑑𝑑=𝑑2𝑠𝑑𝑑2) d) Use your result from part c to calculate the velocity at 𝑑=2𝑠 and 𝑑=6𝑠. e) Compare your results for part b and part d.


1
Expert's answer
2022-02-23T12:35:23-0500
"s(t)=10t+\dfrac{4t^2}{2}, m"

a)


"s(t)=10t+2t^2, 0\le t\le10"


b)


"grad \ s=\dfrac{s_2-s_1}{t_2-t_1}"

"t=2 , s(2)=28"



"grad \ s|_{t=2}=\dfrac{10(2+\Delta t)+2(2+\Delta t)^2-28}{\Delta t}"


"=18+2\Delta t\to 18( m/s)"

"t=6 , s(6)=132"



"grad \ s|_{t=6}=\dfrac{10(6+\Delta t)+2(6+\Delta t)^2-132}{\Delta t}"


"=34+2\Delta t\to 34( m/s)"

c)

i)


"v(t)=\dfrac{ds}{dt}=10+4t, m/s"

ii)


"a(t)=\dfrac{dv}{dt}=\dfrac{d^2s}{dt^2}=4\ m/s^2"


d)


"v(2)=10+4(2)=18(m\/s)"

"v(6)=10+4(6)=34(m\/s)"


e) The results are the same.